Mr. Wade carries 3 more than 3 times as many different magazines as newspapers. He carries 85 different newspapers. How many different magazines does Mr. Wade carry?
1
Expert's answer
2013-04-08T08:38:54-0400
Let m be the number of different magazines and n=85 be the number of different newspapers.Then by the assumption
m = 3 + 3 n
= 3 + 3*85
= 3 + 255
= 288
Thus Mr. Wade carries 288 different magazines.